I need help with calculating Deciles! I have calculated it but my answers seems to be wrong! I do not understand what my mistake is. :(
a. D1 b. D5 c. D9
Class Interval-----Frequency-------Cumulative Frequency:
18-21----4----60
14-17---16----56
10-13---20----40
6-9-----14----20
2-5-----6------6
Summation of frequency: 60

Answer by stanbon(75887)

I need help with calculating Deciles! I have calculated it but my answers seems to be wrong! I do not understand what my mistake is. :(
a. D1 b. D5 c. D9
Class Interval-----Frequency-------Cumulative Frequency:
18-21-------------------4-----------------60
14-17------------------16-----------------56
10-13------------------20-----------------40
6-9--------------------14-----------------20
2-5---------------------6------------------6
Summation of frequency: 60
Decile Formula:: D = LB+[(kN/10 - cfb)/fDk]i
LB is lower limit of Decile group
kN is (Decile number)(sum of freqs)
cfb is sum of frequencies below the Decile group
fDk is frequency of the Decile group
i is the width of the Decile group
D1 =
Decile group is 2-5;LB = 1.5;dN= 1*60;cfb =0;fDk=6;i=4
So D1 = 1.5+[(1*60/10-0)/6]4 = 1.5+1*4 = 5.5
D5 =
Decile group is 10-13:LB=9.5;dN=5*60;cfb=26;fDk =20; i = 4
So D5 = 9.5+[(5*60/10-26)/20]4 = 9.5+(4/20)*4 = 10.3
D9 =
Decile group is 10-13;LB = 9.5;dN=9*60;cfb=26;fDk=20;i = 4
So D9 = 9.5+[(9*60/10-26)/20]4 = 9.5+[28/30]4 = 13.23
-----------------
Cheers,
Stan H.
-----------